### v4

# BadBlocks: Lightweight and Stealthy Backdoor Threat in Text-to-Image Diffusion Models

Abstract
Diffusion models have recently achieved remarkable success in image generation, yet growing evidence shows their vulnerability to backdoor attacks, where adversaries implant covert triggers to manipulate outputs. While existing defenses can detect many such attacks via visual inspection and neural network–based analysis, we identify a more lightweight and stealthy threat, termed BadBlocks. BadBlocks selectively contaminates specific blocks within the UNet architecture while preserving the normal behavior of the remaining components. Compared with prior methods, it requires only about 30% of the computation and 20% of the GPU time, yet achieves high attack success rates with minimal perceptual degradation. Extensive experiments demonstrate that BadBlocks can effectively evade state-of-the-art defenses, particularly attention-based detection frameworks. Ablation studies further reveal that effective backdoor injection does not require fine-tuning the entire network and highlight the critical role of certain layers in backdoor mapping. Overall, BadBlocks substantially lowers the barrier for backdooring large-scale diffusion models, even on consumer-grade GPUs.
Figure 1: We propose BadBlocks, a novel backdoor attack method that fine-tunes the entire UNet by selectively training only specific sampling blocks. By significantly reducing the number of trainable parameters, BadBlocks lowers the computational and GPU resource requirements for the attacker. Moreover, since the remaining blocks are kept entirely frozen, the overall model performance is preserved, resulting in negligible FID degradation.
I Introduction
Today, generative models have become some of the most widely adopted models in the field of artificial intelligence, with diffusion models emerging as one of the most effective approaches. Diffusion models are extensively used to generate high-quality images and videos, and they support a wide range of conditional inputs, such as text prompts, reference images, and ControlNet maps to guide the generation process[diffusion_survey1, diffusion_survey2].
However, recent studies have shown that diffusion models are vulnerable to backdoor attacks in which attackers can manipulate the output of models by embedding covert triggers in the input[trojdiff, baddiffusion]. Backdoor attacks pose significant challenges to the safe development of diffusion models. Once the backdoor is activated by attackers, it can lead to unpredictable and potentially harmful outcomes, such as the generation of malicious content or the manipulation of decisions in downstream classifiers[survey, survey2, survey3]. In addition, due to the stealthy nature of backdoors and the diversity of potential triggers, backdoor attacks are widely regarded as one of the most serious threats to diffusion models. In threat scenarios, attackers upload backdoor-injected models to a public platform (e.g., Hugging Face or Github) while falsely claiming it to be benign. When unsuspecting users download and deploy these models, their output typically appears normal and harmless. However, once the attacker activates the backdoor, the target model begins to generate predefined malicious content[BadT2I].
The powerful generative capabilities of diffusion models, coupled with their vulnerability to backdoor attacks, have raised widespread concerns. With the increasing release of pre-trained models, users frequently download the necessary models from open platforms according to their specific requirements[pretrained:1, pretrained:2]. However, due to the opacity of black-box models, users often focus solely on whether the pre-trained models can accomplish the intended tasks, while overlooking potential backdoor threats[backdoor:1, backdoor:2, backdoor:3]. For example, in classification tasks, attackers can easily manipulate the output by activating backdoors[classback]. In face recognition tasks, when facial images contain trigger patterns, the model may misclassify them as any identity predetermined by the attacker[faceback]. In generative tasks, backdoor attacks often cause the model to produce malicious content, including violent, gory, or otherwise inappropriate imagery, which may further compromise downstream applications[llmback]. In diffusion models, the objective of backdoor attacks is often to generate a specific image or a particular category of images.
Fortunately, successfully executing covert backdoor attacks remains a highly challenging task. On the one hand, various defense frameworks have been developed to counter backdoor attacks in diffusion models. These frameworks typically achieve backdoor detection and trigger inversion by designing specialized neural network architectures and tailored loss functions[eliagh, terd]. On the other hand, attackers often incur significant costs for computing resources and must invest considerable time in fine-tuning the model to inject covert and effective backdoors, which limits the scalability of backdoor attacks.
However, in this paper, we find that the influence of backdoor attacks on the parameters of the diffusion model can be modeled as fine-grained sample blocks. Based on this discovery, we reveal a novel threat that enables attackers to inject backdoors into diffusion models with minimal cost and significantly reduced training time. Notably, BadBlocks exhibits strong stealthiness and effectively mitigates the assimilation phenomenon in attention layers. Our main contributions are summarized as follows:
  • We propose BadBlocks, a lightweight, stealthy, and fine-grained backdoor attack framework that allows attackers to inject backdoors into diffusion models by training only on the most vulnerable sample blocks. Experimental results show that, compared with previous approaches, BadBlocks reduces video memory usage by up to 70% and GPU training time by 80%, and can be executed on standard consumer-grade graphics cards.
  • For the first time, we identify that different layers of neural networks in diffusion models exhibit varying degrees of sensitivity to backdoor attacks. We model the differences among these parameters and highlight the critical role of different components in facilitating backdoor attacks. In addition, we conduct an in-depth analysis of the causes of the ”assimilation phenomenon” in Text-to-Image backdoor and reveal the limitations of backdoor attack detection via defense mechanisms based on attention features.
  • In our ablation study, we conducted a fine-grained analysis of neural network blocks and layers to investigate the role of different components in backdoor attacks. Previous studies have primarily examined the impact of backdoor attacks at the model level. To the best of our knowledge, this is the first work to analyze backdoor vulnerabilities at the layer level, offering a theoretical foundation for future research on the interpretability of backdoor mechanisms.
II Related Works
This section reviews the fundamentals of diffusion models, existing backdoor attacks, and representative defense strategies, and summarizes their limitations.
II-A Fundamental Principles of Diffusion Models
Diffusion models are powerful generative models that learn data distributions from training samples $x$. Denoising Diffusion Probabilistic Models (DDPM) first formalized diffusion-based image generation by modeling a forward noising process $x_{0}\rightarrow x_{t}$ and a reverse denoising process[ddpm]. The forward process follows $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\overline{\alpha}}\epsilon$, while the reverse step is given by $\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\epsilon,\ \epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, where $\epsilon_{\theta}$ is the model prediction. Denoising Diffusion Implicit Models (DDIM) remove the Markovian dependency to accelerate sampling[ddim]. Latent Diffusion Models (LDMs) compress data into a latent space using a VAE to reduce computational cost[stablediffusion, vae]. Stochastic Differential Equations (SDE) unify diffusion processes under continuous formulations and introduce score-based training[sde]. Further, ODE-based schedulers enable fast sampling within 15–30 steps[dpmv1, dpmv3].
II-B Backdoor Attacks in Diffusion Models
Backdoor attacks have emerged as a critical threat to diffusion models[riskindiff]. By fine-tuning models, attackers establish covert neuron-level mappings to force predefined outputs[backdoor:4]. TrojDiff first injected trigger-specific perturbations into the noise space of DDPM and DDIM[trojdiff], while BadDiffusion directly embedded patches into the noise space without trigger functions[baddiffusion]. RickRolling extended attacks to the textual prompt space in LDMs, especially Stable Diffusion[rickroll]. VillanDiffusion unified noise and text attack spaces for broader targets[villan]. Beyond these, Gungnir exploited image style as triggers in image-to-image tasks[gungnir]. Overall, the backdoor attack space is rapidly expanding, posing increasing security risks.
II-C Defense Strategies in Diffusion Models
To counter these threats, various defenses have been proposed. Elijah is the first framework tailored for diffusion models, mitigating attacks such as BadDiff, TrojDiff, and VillanDiffusion[eliagh]. TERD extends Elijah by modeling noise-space attacks as $x_{t}=a(x_{0},t)x_{0}+b(t)\epsilon+c(t)r$, and detects backdoors via KL divergence between benign and inverted noise distributions. For text-based attacks, T2IShield analyzes attention maps and the ”Assimilation Phenomenon” to defend against prompt-triggered backdoors like RickRolling[T2IShield]. Recent defenses increasingly rely on distributional analysis within black-box models rather than handcrafted architectures. However, emerging scenarios such as ControlNet and adapter-based tasks remain underexplored[IP-Adapter].
II-D Limitations
Existing attacks face two main challenges: (1) high computational cost comparable to full model training, and (2) target-specific poisoning often induces the ”assimilation phenomenon,” increasing detectability. From the defense perspective, current methods lack generalization across attack spaces: Elijah and TERD focus on noise-based attacks, whereas T2IShield is limited to prompt-based scenarios.
III Method
In this section, we introduce the BadBlocks threat model, outlining its attack scenarios and background. We analyze the capabilities of both attackers and defenders, then briefly describe the attack method with a theoretical explanation of its feasibility.
III-A Threat Model
The threat model for backdoor attacks on diffusion models is typically defined as a black-box setting from the user’s perspective[bagm, picc]. Depending on the attack method, adversaries can poison the training data or alter the training strategy, ultimately providing backdoored models to the user[dpf]. The threat model in BadBlocks is largely consistent with previous work. However, because the weight distribution and features of the remaining blocks remain unchanged, BadBlocks is more likely to evade detection by defense frameworks, especially which methods based on the global features of the UNet structure. You can find the analysis of the differences in the distribution of model weights in Appendix A and B. Building on existing work[BadNL, DBA], we define the knowledge of the attacker in the BadBlocks setting as follows: (1) The attacker has the authority to manipulate the fine-tuning process of the model, including the dataset, the loss function and the training strategy; (2) The attacker can access the model and provide input data that comply with its specifications. In contrast, a model defender not only has full access to all model parameters and can input arbitrary data to obtain detection samples, but can also perform additional operations such as distillation, pruning, and other modifications about model weights [distillerase, diff-clean].
Figure 2: In infected UNet models, the normalization layers exhibit the most significant weight changes, followed by the ResNet layers and Transformer blocks. We hypothesize that this hierarchy of changes is critical for establishing effective backdoor mappings.
III-B Mitigation Assimilation
The key intuition behind the design of BadBlocks is that the assimilation phenomenon is closely tied to the cross-attention layers across various blocks in the diffusion model. Therefore, to effectively eliminate assimilation, it is crucial to preserve benign weights in as many attention layers as possible. This raises two key questions: (1) Can the impact of backdoor-infected weights be minimized? (2) Are all UNet blocks and neural network layers essential for executing a backdoor attack?
Inspired by previous studies[hm, le], we recognize that different neural network blocks within the UNet architecture of a diffusion model serve distinct roles during the image generation process. Therefore, we attempt to design an algorithm that decouples the neural components associated with backdoor mapping from the UNet and applies fine-grained training to them. Specifically, in the attention-based detection method, given a token sequence of length $L$, each block $b$ generates a cross-attention map $M_{b}^{(t)}$ at any time step $t\in T$:
$M_{b}^{(t)}=\text{CrossAttn}(Q^{(t)},K^{(t)},V^{(t)})\in\mathbb{R}^{L\times D},$ (1)
$T$ is a hyperparameter that determines the number of iterations of sampling performed by the model. We denote the network depth of UNet as $D$, representing all the sampling blocks in the network. By averaging the attention maps throughout the sampling process, we obtain the attention map sample $\hat{M}$ used for backdoor detection:
$\bar{M}_{b}=\frac{1}{|T|}\sum_{t=1}^{T}M_{b}^{(t)},$ (2)
$\hat{M}=\frac{1}{|D|}\sum_{b=1}^{D}\bar{M_{b}}.$ (3)
In the defense evaluation of BadBlocks, we adopt the F-Norm Threshold Truncation method proposed in T2IShield, a lightweight detection approach grounded in the assimilation phenomenon. Specifically, by analyzing the F-Norm distribution, we derive the detection sample F:
$F=\|\hat{M}\|_{F}=\frac{1}{L}\sum_{t=1}^{L}\sqrt{\sum_{i=1}^{D}\sum_{j=1}^{D}(\bar{M_{b}}-\hat{M})^{2}},$ (4)
Finally, a threshold value $\hat{F}\in\mathbb{R}^{1}$ is applied to determine whether the sample contains a backdoor trigger:
[…]
However, in BadBlocks, the attention maps $M_{b}$ generated by most blocks appear to be entirely normal, as their parameters remain unchanged. Only the injected sampled blocks that are targeted by the attack exhibit abnormal attention patterns, and these anomalies can be easily masked when averaged with the attention maps from benign blocks. This allows the remaining benign cross-attention layers to suppress the assimilation effect induced by the backdoor. Ultimately, through controlled experiments, we successfully isolated the network blocks essential for executing backdoor attacks in diffusion models, specifically, network blocks in the upsampling stage and its critical subcomponents: the the resnets[resnet], the transformer blocks[trans] and the normalization layers[groupnorm] among them.
TABLE I: We evaluated the minimum computational overhead of BadBlocks for 20,000 samples on an NVIDIA A40 GPU. Compared to existing methods, BadBlocks introduces significantly less impact on the target model. Moreover, its low dependency on attacker-side hardware substantially lowers the barrier to executing backdoor attacks.
Method Infected Components Affected Parameters Necessary GPU Memory GPU Time (1 epoch)
BadT2I UNet 859M/1066M 18529MB 7.5 hour
VillanDiffusion UNet & Text Encoder 1066M/1066M 21425MB 9.5 hour
RickRolling Text Encoder 123M/1066M 6971MB 3.2 hour
BadBlocks(Ours) UpSample Blocks 18M/1066M 5589MB 1.4 hour
Figure 3: Our findings indicate that the final upsampling block plays a critical role in enabling backdoor mapping, while also demonstrating that not all model parameters are essential for executing backdoor attacks.
III-C Approaches of BadBlocks
The training of diffusion models typically involves a model $M$ parameterized by weights $\theta$, which learns to approximate the data distribution from the original dataset $x$. The trained model is capable of generating image $x^{\prime}$ from noisy inputs $\epsilon$ conditioned on additional information $a$, which can be formally expressed as:
$x^{\prime}=M_{\theta}(\epsilon,a,t),\epsilon\sim{N(0,I)},$ (6)
where $t\in{T}$ is the variance of timestep.
We adopt the definition of the input space $S_{input}$ from Gungnir[gungnir], where the additional conditions space $a\subseteq{A_{cond}}$ is introduced to constrain the stochasticity of the diffusion process. The complete composition of the additional space $A_{cond}$ is:
$A_{cond}=\{prompts,images,controlnet,$...\}. (7)
Therefore, the complete input space $S_{input}$ for the diffusion process can be defined as follows:
$S_{input}=\{(\epsilon,a)|\epsilon\sim{N(0,I)},a\subseteq{A_{cond}}\}.$ (8)
In BadBlocks, we define a hyperparameter $i\geq 0$ to represent the number of blocks involved in backdoor training, where $i$ is strictly less than the total number of upsample blocks. Importantly, during training, we froze nearly all parameters of the model $M_{\theta}$, performing gradient updates only on the critical weight $\theta^{*}_{i}$, including the ResNet layers, Transformer blocks, and normalization layers. Figure reffig:fig4 illustrates the specific parameters within the sampled blocks targeted by BadBlocks that require training.
Algorithm 1 Overall BadBlocks Training Procedure
0: Target model $M$, Clean dataset $\mathbf{D_{c}}$, Secret trigger $\mathbf{s}$, Backdoor target $\mathbf{i_{t}}$, Number of BadBlocks $i$  Model parameters $\theta$, Poisoning rate $r$, Learning rate $\eta$;
0: Backdoored model $M^{{}^{\prime}}_{\theta^{*}}$;
[…]
5:while remaining epochs do
6:  $x_{0}\sim$ Uniform $\mathbf{D_{p}}$;
7:  if backdoortrainingbackdoor\ training then
8:   $t\sim$ Uniform(1,,T)({1,...,T});
9:   $\mathrm{d}\mathbf{x}_{t}=f(t,\mathbf{x}_{t})\,\mathrm{d}t+g(t)\,\mathrm{d}\mathbf{w}_{t}$; # Add random noise
10:   $\mathcal{L}=\mathbb{E}_{x_{0},s,t}\left[\left\|\epsilon-\epsilon_{\theta^{*}_{i}}(x_{t},a_{b},t)\right\|^{2}_{2}\right]$;
11:  end if
12:  Update $\theta^{*}_{i}$ using optimizer step on loss $\mathcal{L}$
13:end while
14:return $M_{\theta^{*}}^{{}^{\prime}}$;  # Return the backdoored model
Owing to the universality of BadBlocks, its additional condition space in backdoor attacks, denoted as $a_{b}$, closely approximates the entire benign input space $A$. This indicates that BadBlocks can accommodate nearly any existing trigger dimension. Based on this input space, we define the final training loss function for the BadBlocks backdoor as follows:
[…]
Figure 4: One key advantage of BadBlocks is that it does not rely on the modification of the loss function, enabling broad compatibility with existing backdoor methods. It consistently produces high-quality images across various triggers with minimal degradation.
Algorithm.1 presents the fine-tuning process of BadBlocks, where the training procedure is carried out by freezing the majority of weights. Furthermore, to quantify the weight changes introduced by backdoor attacks, we compute the absolute mean difference $\mathbf{d}$ between the weights of the benign and backdoor model. Every $d_{i}$ between each pair of benign and backdoor weights can be expressed as:
$d_{i}=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\left|\theta_{i,j}^{(benign)}-\theta_{i,j}^{(backdoor)}\right|,$ (10)
$\mathbf{d}=[d_{1},d_{2},\dots,d_{k}],$ (11)
where $k$ is the total number of parameter tensors in the block. In the experimental section, we leverage this differential analysis to identify the parameters that play a pivotal role in establishing backdoor mappings within neural networks, thereby providing critical evidence to support our findings. Figure 10 visualizes the differences in weight changes across components affected by BadBlocks.
IV Experiments
In this section, we present the evaluation results of BadBlocks in the text-to-image task, along with its corresponding attack settings. Figure 4 shows the attack process of BadBlocks. Furthermore, we highlight the advantages of our method in terms of training efficiency and resource consumption while also acknowledging its limitations.
IV-A Attack Configuration
In the experimental section, we adopted standard settings for the backdoor attack. For the text-to-image task, we used three different characters as the backdoor trigger, including invisible Unicode, specific phrase, and lowercase greek letter. As attack targets, we selected three mainstream diffusion models: Stable Diffusion v1.5, Stable Diffusion v2.1 and Realistic Vision v4.0, which were trained on the MS-COCO-Caption dataset[coco]. To assess the quality of the generated images, we use the FID score as the primary metric[FID]. Furthermore, we employ LPIPS and SSIM to evaluate the attack success rate (ASR)[LPIPS, ssim]. All training is conducted for 5 epochs, with a learning rate of 1e41e-4 and a batch size of 8. All training and evaluation processes were conducted on an NVIDIA A40 GPU, but this does not mean that professional computing graphics cards are necessary for BadBlocks. Our experimental results demonstrate that consumer-grade GPUs, such as the NVIDIA RTX 3060 (12GB) and RTX 4070Ti SUPER (16GB), can also perform efficient and successful backdoor attacks within a short time frame. Table.I presents a comparison between BadBlocks and existing approaches in terms of parameter modification scope and GPU resource consumption.
IV-B Attack Results
Table II reports the attack performance of BadBlocks on three baseline models. The results show that establishing an effective backdoor mapping requires updating only a small subset of key parameters, indicating strong redundancy in diffusion models that can be exploited by adversaries. On Stable Diffusion v1.5, BadBlocks incurs only 4.3% FID degradation, while the UNet-based method suffers a loss of 73.4% under the same poisoning rate. Similar trends are observed on the other two baselines, demonstrating the generality of our method. These results validate the feasibility of fine-grained backdoor training via architectural decoupling, suggesting that infecting only critical components is sufficient to achieve high attack success. Furthermore, by substantially reducing parameter modifications, BadBlocks lowers computational cost, shortens training time, and alleviates GPU memory requirements. Owing to its independence from specific loss manipulations or additional regularization, BadBlocks is highly compatible with existing attack strategies. As shown in Appendix.E, integrating BadBlocks with mainstream attacks significantly reduces resource consumption without sacrificing attack success or stealthiness. Additional results under different schedulers are provided in Appendix.C.
TABLE II: We evaluated the performance of BadBlocks across three different baseline models and found that it not only executed backdoor attacks effectively, but also exhibited strong stealthiness.
Models ASR $\uparrow$ FID Score $\downarrow$ Target Similarity
[…]
Figure 5: BadBlocks maintains or improves generation quality (lower FID) after backdoor injection, with minimal usability loss compared to other UNet-based attacks. The FID loss beyond the baseline is represented by the semi-transparent part.
V Ablation Study
In this section, we investigate the impact of different configurations on the attack performance, including ablations of key components and the different number of infected blocks.
V-A Impact of Crucial Components
To explore component-level roles in backdoor attacks and reduce the attack surface, we conducted ablation studies on the ResNet layer, Transformer block, and normalization layer within the sampling block. Results show that single-component fine-tuning fails to inject a backdoor, while jointly attacking the normalization and Transformer blocks enables backdoor injection at the cost of degraded image quality. As shown in Figure 6, although the ResNet layer is not directly involved in backdoor expression, it is essential for preserving normal generation quality.
Figure 6: The activation of the backdoor process cannot be achieved without the participation of key network components. A single component is insufficient for backdoor mapping, we found effective attacks require at least the ResNet and Transformer blocks in upsampling blocks.
V-B Different Number of Infected Blocks
In general, the upsampling stage of a diffusion model consists of multiple sequentially connected upsampling blocks. We conducted ablation studies by varying the number of infected sample blocks used for fine-tuning. The experimental results reveal that increasing the number of fine-tuned blocks does not necessarily enhance backdoor performance. In contrast, an excessive number of infected blocks may hinder effective backdoor mapping, suppress backdoor activation, and induce excessive assimilation effects. Figure 7 shows model performance with backdoor injection across different numbers of injected blocks.
Figure 7: BadBlocks performance correlates with the number of infected blocks. It is evident that the attack performance is not positively correlated with the number of affected parameters, suggesting that certain components within the model are particularly sensitive to backdoor injection.
VI Conclusion
In this work, we uncover a novel threat, BadBlocks, which enables adversaries to implant stealthy backdoors by fine-tuning only a subset of neural network blocks within diffusion models. Unlike existing approaches, BadBlocks effectively circumvents the assimilation phenomenon while offering key advantages such as reduced computational overhead, faster training, and minimal disruption to model parameters. We advocate for continued investigation into the structural mechanisms and critical parameter regions that facilitate backdoor injection. Furthermore, we emphasize the need for robust and generalizable defense strategies capable of countering attacks like BadBlocks, thereby safeguarding the integrity and reliability of image generation systems.
Appendix A Technical Appendix
To assess the concealment of BadBlocks, we designed three additional experiments, utilizing L-norm and weight statistics evaluate the impact of BadBlocks on the weights of the base model. The results of these experiments demonstrate that BadBlocks is a highly concealed and difficult-to-detect threat, with a much smaller impact on the weights of the base model compared to previous methods. In addition, we evaluated the performance of BadBlocks across different diffusion schedulers. The sampling results showed that BadBlocks is a generalizable backdoor attack method.
A-A A.Evaluating BadBlocks via L-norms
The L-norm is commonly used in machine learning to measure the distance between two vectors.The method of calculating the degree of dispersion varies depending on the measurement parameter $p$. The L-P norm can be expressed as:
$\|\mathbf{x}\|_{p}=\left(\sum_{i=1}^{n}|x_{i}|^{p}\right)^{\frac{1}{p}},$ (12)
When $p=2$, it is referred to as the L-2 norm, or the Euclidean norm, and is used to calculate the Euclidean distance from a vector to the origin. The L-2 norm is commonly used to represent error or loss functions, particularly when measuring the difference between vectors. We use this function to measure the difference between the infected and pure weights under different attack configurations, which can be expressed as:
$\|Diff\|_{2}=\sqrt{\sum_{i=1}^{n}(W_{benign}^{i}-W_{backdoor}^{i})^{2}},$ (13)
where $W^{i}$ represents the weight block $i$ in the model and $n$ is the total number of weight blocks. We compared BadBlocks with VillanDiffusion and RickRolling. The experimental results show that the L-2 norm distance between BadBlocks and the base model is significantly lower than that of other methods, which indicates that the backdoor mapping created by BadBlocks exhibits greater stability, making it harder to detect by defense methods based on parameter analysis. Figure.8 shows the differences in the L-2 norm among different attack methods.
Figure 8: BadBlocks exhibits a significant difference from other UNet-based methods in L-2 norm measurements, indicating that it is closer to the benign model in terms of Euclidean distance.
A-B B.The variation of parameter distribution
Previous studies have consistently analyzed backdoor distributions and researchers generally believe that covert backdoor attack methods should closely match a benign distribution. Therefore, we analyzed the absolute differences between the weights of the infected and base models in different attack methods(BadBlocks and UNet-based) and performed a statistical analysis of the results. In Figure.9, we computed the average mean and standard deviation of these weight differences:
[…]
$std^{2}=\frac{1}{N}\sum_{i=1}^{N}\left(\left|W^{(i)}_{\text{benign}}-W^{(i)}_{\text{backdoor}}\right|-mean\right)^{2}$ (15)
Experimental results show that, compared to UNet-based methods, BadBlocks exhibits significantly smaller numerical differences in model weights. This indicates that the weights modified by BadBlocks are closer to the benign model in vector space, and also suggests that the infected model has greater resistance to parameter-clipping-based defense algorithms.
Figure 9: Statistical results indicate that the weight difference of BadBlocks is significantly smaller than that of UNet-based methods.
A-C C.Performance under different schedulers
We conducted a comprehensive evaluation of BadBlocks across four widely used diffusion sampling methods—DDIM, LSMD, DPM, and PNDM—to assess its generalizability and robustness under different generation frameworks. These schedulers vary in their sampling strategies: DDIM provides a deterministic approximation of the reverse diffusion process, LSMD utilizes a continuous-time stochastic differential equation (SDE) formulation, DPM offers high-speed generation with reduced discretization error, and PNDM combines multiple-step prediction with momentum-based refinement.
The experimental results demonstrate that BadBlocks consistently maintains high attack performance across all tested schedulers, achieving an attack success rate approaching 100%, while introducing only minimal degradation in image quality, as measured by the Fréchet Inception Distance (FID), under the optimal Poisoned Rate (PR). The detailed evaluation outcomes are illustrated in Figures 3, highlighting the stable and effective performance of BadBlocks under each sampling method.
Figure 10: Attack Performance under different schedulers.
A-D D. Visualized Attack Results
Figure 11: Other target images.
A-E E.Optimize Existing work with BadBlocks
One key advantage of BadBlocks is its compatibility with most mainstream attack methods, allowing them to retain their original attack effectiveness while gaining the enhanced stealth provided by BadBlocks. In this section, we select three representative attack methods—RickRolling, BadDiffusion, and VillanDiffusion. We apply BadBlocks to enhance their stealth and optimize overall attack performance.
A-E1 RickRolling
RickRolling is the first work to extend the attack surface to prompt by using special characters as triggers. These characters are encoded by the text encoder into attack vectors, which subsequently trigger backdoor mapping through the Cross-Attention layer. After applying BadBlocks to lock the non-critical weights, we found that RickRolling was still able to carry out effective backdoor attacks. Not only that, using any prompt-based trigger can maintain good concealment and attack performance.
In this section, Stable Diffusion v1.5 is used as the baseline model, and experiments are conducted on the MSCOCO dataset with a minimum batch size of 1. All experiments were performed on an NVIDIA A40 GPU.
Trigger FID ASR FTT 1 Epoch
o(U+03BF)
31.83
19.62
98.78%
99.01%
97.32%
9.42%
3.2 hour
1.2 hour
o(U+043E)
28.47
17.90
99.12%
98.65%
95.84%
10.25%
3.0 hour
1.2 hour
L(U+2113)
27.68
19.95
97.98%
98.32%
94.21%
8.87%
3.1 hour
1.3 hour
”zebra apple”
26.21
18.13
98.55%
97.82%
96.77%
9.98%
3.1 hour
1.3 hour
”trigger”
27.35
17.88
99.03%
98.47%
97.10%
10.55%
3.2 hour
1.3 hour
TABLE III: It is evident that integrating BadBlocks can significantly enhance attack efficiency and generation quality, while further improving resistance to attention-based detection methods.
In Table.III, after applying BadBlocks, RickRolling’s image generation quality increased by up to 38.3%, reduced BDR by up to 90.6%, and cut GPU time by up to 62.5%. This is sufficient to demonstrate the advantage of BadBlocks in terms of attack performance.
A-E2 BadDiffusion
BadDiffusion is a well-established attack method within the image input space that activates the backdoor mapping by embedding a patch trigger into the image. Building upon BadDiffusion, we have developed an algorithm based on BadBlocks specifically for image-to-image tasks. Details of the algorithm can be found in Algorithm.2. We will discuss the attack performance in Table.IV.
Algorithm 2 BadBlocks-Based BadDiffusion for Image-to-Image Tasks
0: Target model $M$; Clean image-to-image dataset $\mathbf{D_{c}}$; Visual trigger patch $\mathbf{s}$; Backdoor target image $\mathbf{i_{t}}$; Number of BadBlocks $i$; Model parameters $\theta$; Poisoning rate $r$; Learning rate $\eta$;
0: Backdoored model $M^{{}^{\prime}}_{\theta^{*}}$;
[…]
6:while not converged do
7:  $(x_{0},prompt)\sim\mathbf{D}$;  # Sample a training pair
8:  $t\sim\text{Uniform}(1,\ldots,T)$;  # Random timestep
9:  $\mathbf{x}_{t}\leftarrow\text{AddNoise}(x_{0},t)$;  # Diffusion forward process
10:  $\mathcal{L}=\mathbb{E}_{x_{0},s,t}\left[\left\|\epsilon-\epsilon_{\theta^{*}_{i}}(\mathbf{x}_{0},s_{b},t)\right\|_{2}^{2}\right]$;
11:  Update $\theta^{*}_{i}$ using optimizer and loss $\mathcal{L}$;
12:end while
13:return $M^{{}^{\prime}}_{\theta^{*}}$;  # Return the Model
Trigger FID ASR FTT Memory
Square
25.25
18.73
97.45%
98.19%
3.4 hour
1.5 hour
18254MB
5601MB
Stop Sign
23.10
17.42
98.91%
98.33%
3.2 hour
1.5 hour
18260MB
5612MB
Glasses
25.05
18.96
97.83%
97.95%
3.3 hour
1.5 hour
18257MB
5601MB
TABLE IV: Since BadDiffusion relies on patch-based triggers, BadBlocks cannot reduce the visual prominence of these patches, and thus existing defense frameworks targeting BadDiffusion may remain effective. However, this does not negate the reduced FID loss and lower resource consumption achieved by integrating BadBlocks.
A-E3 VillanDiffusion
VillanDiffusion is a unified backdoor attack framework for diffusion models, with an algorithmic structure similar to that of RickRolling and BadDiffusion. The key distinction is that VillanDiffusion offers a detailed performance evaluation specifically for latent diffusion models (LDMs) under a stochastic differential equation (SDE) framework.
Therefore, in this part, we present detailed results of BadBlocks under various sampling methods, including both ODE-based and SDE-based samplers such as DDIM, DDPM, DPM-Solver, and PNDM. As shown in Table.V, a comparison of attack performance reveals that nearly all schedulers exhibit performance improvements when integrated with BadBlocks.
Scheduler FID ASR 1 Epoch Memory
DDPM
30.25
18.73
97.45%
98.19%
3.4 hour
1.5 hour
18254MB
5601MB
DDIM
28.32
17.41
98.60%
98.33%
3.1 hour
1.5 hour
18260MB
5612MB
DDIM
27.45
18.10
98.10%
97.85%
3.2 hour
1.4 hour
18257MB
5603MB
DPM-o1
26.91
17.35
98.74%
98.42%
2.9 hour
1.5 hour
18102MB
5567MB
DPM-o2
27.13
16.88
99.01%
98.67%
2.7 hour
1.6 hour
18095MB
5542MB
PNDM
29.52
18.03
98.24%
98.20%
3.0 hour
1.4 hour
18248MB
5593MB
TABLE V: Even under different schedulers, BadBlocks consistently activates backdoors and achieves exceptionally high attack success rates . Obviously, the random noise introduced by SDE-based sampling fails to disrupt the backdoor mapping of BadBlocks, further highlighting its robustness as a persistent threat.
Experimental results demonstrate that BadBlocks significantly enhances the efficiency of backdoor attacks on diffusion models across nearly all schedulers, while also substantially reducing computational overhead.